OpenStudy (anonymous):

A rock is dropped form the top of a 400 feet high cliff. Assuming that gravity is the only force acting on the rock: A) What is the initial velocity of the rock? B) Determine the formula, v(t), for the velocity of the object at time t. C) What is the initial height of the rock? D) Determine the formula, s(t), for the distance the rock is above the ground at time t. E) At what time t does the rock hit the ground? F) What is the terminal velocity of the rock (i.e. the velocity of the rock when it hits the ground)?

OpenStudy (anonymous):

a)\[V²=U²+2as\] consider that U(initial speed = 0) V²=2 x 9.8m/s x 121.92m ----(400 ft) \[V=\sqrt{2389.632}\] V=48.88m/s

OpenStudy (anonymous):

B)\[V=s/t\] D) V=U+at s=1/2(u+v)t s=ut+1/2at² E) from any equation given in D, u will get approximate 4.8s = t